OpenStudy (youngsuperman):

Francine and Cheryl received equal scores on a test made up of multiple choice questions and an essay. Francine got 34 multiple choice questions correct and received 14 points for her essay. Cheryl got 30 multiple choice questions correct and received 22 points for her essay. How many points was each multiple choice question worth?

OpenStudy (superflytyguy):

34x+14 = 30x+22 solve for x 34(2) = 68 30(2) = 60 68 + 14 = 82 60 + 22 = 82

OpenStudy (superflytyguy):

Each question is worth 2 points each

OpenStudy (youngsuperman):

okay so do u know it

OpenStudy (youngsuperman):

@Townsend5643

OpenStudy (superflytyguy):

Yeah I said each question is 2 points each

OpenStudy (youngsuperman):

yo help with two more

OpenStudy (superflytyguy):

K